OpenStudy (anonymous):

What is the perpindicular line of (-5,2) y=1/2x-3

OpenStudy (anonymous):

-2x-1

OpenStudy (anonymous):

To find a perpendicular line to an equation, inverse the slope of the equation. So, y=2x-3

OpenStudy (anonymous):

How did you figure that?

OpenStudy (anonymous):

the recipricol of 1/2 is -2

OpenStudy (anonymous):

so now you have 2=-2(5)+b

OpenStudy (anonymous):

and solve for b

OpenStudy (anonymous):

nvm its actually y=-2x-8

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

im having troubles

OpenStudy (anonymous):

2=-2(-5)+b

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

how did you get -2 from 1/2

OpenStudy (anonymous):

shouldn't it be a postive 2 not a negative

OpenStudy (anonymous):

(2,6), y=-1/4x+3

OpenStudy (anonymous):

its is opposite reciporical

OpenStudy (anonymous):

not just opposite

OpenStudy (anonymous):

1/2 becomes 2/1 and then becomes negative

OpenStudy (anonymous):

so would i just do : 6=2/1(2)+b

OpenStudy (anonymous):

Then i would have 6=4+b then -4 on both sides and 6-4=2 so b =2????

OpenStudy (anonymous):

or did i do that totally wrong...

OpenStudy (anonymous):

uh oh i lost what i was typing

OpenStudy (anonymous):

a perpendicular slope to 1/2 is -2

OpenStudy (anonymous):

y=-2x+b

OpenStudy (anonymous):

then u plug in the coordinates from above

OpenStudy (anonymous):

and solve for b

OpenStudy (anonymous):

ok so then would i have 6=-2(2)+b

OpenStudy (anonymous):

2=-2(-5)+b 2=10+b -8=b

OpenStudy (anonymous):

ok thanks for the help i think i get this now

OpenStudy (anonymous):

then you remove the x and y coordinates and you have y=-2x-8